# An improvement toward global boundedness in a fully parabolic chemotaxis with singular sensitivity in any dimension

### v1

@@ 4 Proof of the main result
$M(T)\leq c_{3}\left(1+M(T)^{\frac{r}{p_{0}}}\right)\qquad\text{for any }T<T_{\rm max},$
for some $c_{3}>0$ independent of $T$. Since $\frac{r}{p_{0}}<1$, this entails that $M(T)\leq c_{4}$, for some $c_{4}>0$ independent of $T$, which further deduces that $u\in L^{\infty}\left((0,T_{\rm max});L^{p}(\Omega)\right)$. ∎
The following lemma enables us to derive $L^{\infty}(\Omega\times(0,T_{\rm max}))$ bounds for the solution under the assumption that $u\in L^{\infty}((0,T_{\rm max});L^{\frac{n}{2}+}(\Omega))$
@@
Thanks to a uniform lower bound for $v$ as established Lemma 2.3, we follow the standard heat semi-group technique (see Lemma 3.4 in [21]) with the assistance of Lemma 2.4(ii) to prove that $u\in L^{\infty}\left((0,T_{\rm max});L^{\infty}(\Omega)\right)$. . ∎
We are now in position to prove our main result.

### v2

@@ 4 Proof of the main result
$M(T)\leq c_{3}\left(1+M(T)^{\frac{r}{p_{0}}}\right)\hskip 18.49988pt\text{for any }T<T_{\rm max},$
for some $c_{3}>0$ independent of $T$. Since $\frac{r}{p_{0}}<1$, this entails that $M(T)\leq c_{4}$, for some $c_{4}>0$ independent of $T$, which further implies that $u\in L^{\infty}\left((0,T_{\rm max});L^{p}(\Omega)\right)$. ∎
The following lemma enables us to derive $L^{\infty}(\Omega\times(0,T_{\rm max}))$ bounds for the solution under the assumption that $u\in L^{\infty}((0,T_{\rm max});L^{p}(\Omega))$ for some $p>\frac{n}{2}$.
@@
Thanks to a uniform lower bound for $v$ as established Lemma 2.3, we follow the standard heat semi-group technique (see Lemma 3.4 in [27]) with the assistance of Lemma 2.4(ii) to prove that $u\in L^{\infty}\left((0,T_{\rm max});L^{\infty}(\Omega)\right)$. ∎
We are now in a position to prove our main result.
@@
Acknowledgment
The author gratefully thanks the anonymous referees for their valuable comments and suggestions. This work was supported by the Hangzhou Postdoctoral Research Grant (No. 207010136582503) and National Postdoctoral Talent Program.